OpenStudy (anonymous):

simplify completly 25w^2/10x^3 divided by 30w^2/5w

OpenStudy (anonymous):

\[\frac{25w^2}{10x^3} \div \frac{30w^2}{5w}=\frac{25w^2}{10x^3} *\frac{5w}{30w^2}\]\[=\frac{5w}{12x^3}\]

OpenStudy (anonymous):

answers include.. a. 12/5w^2 b. 5w^2/12 c. 15w^4 d. 1/15w^4

OpenStudy (phi):

You must have typed the question wrong. There is no x in any of the answers

OpenStudy (anonymous):

your right.. i meant 10w^3

OpenStudy (anonymous):

i still don't see the answer

OpenStudy (phi):

I also suspect one of your answers is typed in wrong. the answer is \[\frac{5}{12w^2}\]